OpenStudy (anonymous):

Find the area of the shaded region y=(x+2)^(1/2) y=(1/x+1) x= 0, 2

OpenStudy (anonymous):

You first have to know which one is on top

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Looking at the graph, we know that (x+2)^1/2 is on top

OpenStudy (anonymous):

So, Top - Bottom \[\int\limits_{0}^{2}(x+2)^{1/2}-(1/(x+1))\]

OpenStudy (anonymous):

dx

OpenStudy (anonymous):

\[(X+1)^-1\]

OpenStudy (anonymous):

that is what (1/x+1) can be rewritten as right?

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

so now at this point I find anti derivative of the integral right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but when I do i get some weird decimal as an answer, and not the original answer

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\sqrt{x+2} -\int\limits_{0}^{2}(x+1)^{-1}\] u=(x+2) du = dx \[\int\limits_{}^{}\sqrt{u} du = 2/3 u^{3/2}\]

OpenStudy (anonymous):

\[2/3 (x+2)^{3/2}\]

OpenStudy (anonymous):

\[u=x+1\] du=dx \[\int\limits_{}^{}(1/u )du \]

OpenStudy (anonymous):

\[\ln (x+1)\]

OpenStudy (anonymous):

2/3(4)^(3/2)- ln(2)

OpenStudy (anonymous):

whoa we were never introduced into using ln(x+1) as a integral

OpenStudy (anonymous):

Integral of 1/x is ln[x]

OpenStudy (anonymous):

I just received a email from a buddy of mine and end up using this integral in the beginning *next post*

OpenStudy (anonymous):

\[\int\limits_{0}^{6} 2x - (x^2-4x)dx\]

OpenStudy (anonymous):

now i am completely lost on where he gotten that from

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

I mean it would seem to correct to use this approach the way that we are doing the integral, which seems reasonable, but the fact that rewrote the integral into something different confuses me

OpenStudy (anonymous):

I have no idea where this integrand came from

OpenStudy (anonymous):

Can the integral be rewritten in order to get a whole number as a answer, or it impossible to do so with only knowing Calculus 1?

OpenStudy (anonymous):

you get same number no matter how you set up the integral. Sometime they are set up one way or another to make it easier to integrate

OpenStudy (anonymous):

Forgot that I can use U substitution in order to use these integrals so I can figure this out a lot faster. So, I just use this approach and ask my buddy where did he even think about using a different integral.

OpenStudy (anonymous):

Thanks for the help though. I appreciated it